### v2

# Improving Outdoor Multi-cell Fingerprinting-based Positioning via Mobile Data Augmentation

Abstract
Accurate outdoor positioning in cellular networks is hindered by sparse, heterogeneous measurement collections and the high cost of exhaustive site surveys. This paper introduces a lightweight, modular mobile data augmentation framework designed to enhance multi-cell fingerprinting-based positioning using operator-collected minimization of drive test (MDT) records. The proposed approach decouples spatial and radio-feature synthesis: kernel density estimation (KDE) models the empirical spatial distribution to generate geographically coherent synthetic locations, while a k-nearest-neighbor (KNN)-based block produces augmented per-cell radio fingerprints. The architecture is intentionally training-free, interpretable, and suitable for distributed or on-premise operator deployments, supporting privacy-aware workflows. We both validate each augmentation module independently and assess its end-to-end impact on fingerprinting-based positioning using a real-world MDT dataset provided by an Italian mobile network operator across diverse urban and peri-urban scenarios. Results show that the proposed KDE-KNN augmentation consistently improves positioning performance with respect to state-of-the-art approaches, reducing the median positioning error by up to 30% in the most sparsely sampled or structurally complex regions. We also observe region-dependent saturation effects, which emerge most rapidly in scenarios with high user density where the information gain from additional synthetic samples quickly diminishes. Overall, the framework offers a practical, low-complexity path to enhance operator positioning services using existing mobile data traces.
Index Terms:
Mobile Data Augmentation, Fingerprinting, Outdoor Positioning, Minimization of Drive Test (MDT).
I Introduction
Recent years have seen a growing demand for accurate and reliable positioning services in dense urban areas and indoor environments [47]. While satellite-based systems like the global positioning system (GPS) remain the gold standard for global positioning, their utility in operational mobile networks is frequently constrained by practical factors, as they remain power-intensive components [19]. Consequently, mobile operating systems and user policies often disable or duty-cycle satellite tracking to conserve battery life. In these scenarios, multicell fingerprint-based positioning has emerged as a promising approach due to its robustness in non-line-of-sight conditions. However, statistical insufficiency of training data is one of the core limitations of fingerprinting. It manifests in coarse positioning granularity, especially in cell‐dense regions where small geography can exhibit rapid changes in received signal levels, affecting positioning algorithms’ performance. Moreover, collecting exhaustive user traces raises privacy concerns, as these measurements may encode sensitive movement patterns [37]. These challenges motivate the generation of synthetic mobile measurement data that can augment real traces, filling spatial gaps in the fingerprint database while mitigating the need for additional data collection campaigns.
To improve positioning performance in cellular networks, we introduce a mobile data augmentation framework that synthesizes large volumes of realistic radio measurements. Our proposed computationally-light mobile data generator combines kernel density estimation (KDE) for estimating spatial components with non‐parametric regression leveraging k‐nearest‐neighbor (KNN) for multi‐cell reference signal received power (RSRP) features generation. This methodology aims to produce dense, geographically coherent synthetic mobile data that mimics the statistics of real-world measurements. Although this framework is designed and tested end-to-end for fingerprinting-based positioning, the augmented measurements it generates may also benefit other tasks, including radio-environment map reconstruction, training and validation of learning-based models for coverage/capacity estimation and link-quality prediction, data-efficient handover and beam-management policy learning, and anomaly/fault detection in network monitoring [5].
In this work, by leveraging real-world, large-scale minimization of drive test (MDT) data provided by an Italian mobile network operator (MNO), we systematically investigate the positioning performance gains introduced by our proposed synthetic data augmentation pipeline, quantifying improvements in positioning accuracy and robustness across diverse urban scenarios.
I-A Contributions and Paper Organization
This work proposes a novel data augmentation framework specifically designed to enhance fingerprinting-based positioning accuracy in outdoor cellular environments. Our key contributions are summarized as follows:
  • Unlike previous studies that focused primarily on synthetic data augmentation to improve positioning in controlled or indoor scenarios, this work addresses the practical challenges inherent to operational outdoor environments. We validate the proposed framework on real MDT data provided by an Italian MNO in four large-scale settings, demonstrating its capability to handle sparse spatial sampling, heterogeneous and partial physical cell identity (PCI) visibility, measurement noise, and strong temporal/spatial variability typical of user mobility and wide-area propagation effects.
  • The proposed framework introduces a hybrid and modular model that decouples the generation of user locations and radio features, aiming for a lightweight and interpretable architecture. The user spatial distribution is modeled using KDE, while PCI-specific radio fingerprints are generated through KNN.
  • We empirically validate each module of the augmentation pipeline in isolation. For the spatial augmentation stage, we compare KDE against parametric and learned alternatives ( Gaussian mixture model (GMM), generative adversarial network (GAN), and normalizing flow (NF)) using a multivariate two-sample Kolmogorov-Smirnov (KS) test to quantify distributional agreement with real MDT locations. For the radio augmentation stage, we benchmark the proposed KNN approach against random forest (RF), Gaussian process regression (GPR) with both squared exponential (SE) and rational quadratic (RQ) kernels, multilayer perceptron (MLP), conditional generative adversarial network (CGAN), rectified flow (ReFlow) (diffusion model), and NF. We assess prediction quality in terms of median absolute error (MedAE) on per-PCI RSRP values.
  • We assess the end-to-end impact of augmentation on multi-cell fingerprinting-based positioning by comparing the proposed KDE-KNN pipeline against representative two-stage baselines (KDE-RF, KDE-GPR) and advanced deep learning architectures (KDE-MLP, KDE-CGAN, KDE-ReFlow, and KDE-NF) by means of average positioning errors over multiple tests.
  • We quantify how positioning performance scales with the amount of synthetic data. By sweeping augmentation rates and performing pairwise statistical comparisons over positioning errors obtained by our architecture, we identify region-dependent saturation points: beyond a scenario-specific augmentation level, additional synthetic samples yield diminishing returns and, in some cases, may not improve.
  • To facilitate reproducibility and support further research in the field, we make the source code of the proposed framework publicly available. The repository can be accessed at: https://github.com/wilabcnit/Mobile-Data-Augmentation-Positioning.
The remainder of the paper is structured as follows. Sec. II provides the background on fingerprinting and mobile data collection. Sec. III reviews the related work. Sec. IV introduces the proposed mobile data augmentation pipeline, describing the modular components for the spatial and radio features generation. Sec. V discusses both the multi-cell fingerprinting-based positioning problem and introduces the four real-world MDT-based reference scenarios Sec. VI presents the numerical results in terms of (i) independent performance of each stage of the modular framework, (ii) performance benefits of augmentation to fingerprinting-based positioning in the considered scenarios, and (iii) analysis on performance scaling with respect to the amount of synthetic data generated. Finally, Sec. VII concludes the paper.
II Background
II-A Fingerprinting in Cellular Networks
Fingerprint-based positioning leverages existing cellular infrastructure and has become a standard for robust positioning in dense urban and indoor environments [16, 13]. The 3rd generation partnership project (3GPP) incorporated fingerprinting-based positioning into LTE networks under Release 12 [1]. This approach estimates a device’s position by comparing its observed radio measurements, such as RSRP, to a pre-constructed database of location-tagged fingerprints. By leveraging existing cellular infrastructure, fingerprinting provides robust performance where signal blockage and multipath propagation often degrade the accuracy of conventional techniques. However, achieving high positioning accuracy through multi-cell fingerprinting requires large volumes of spatially and radio‐feature‐rich data in order to capture fine‐grained variations across coverage areas; in practice, collecting such dense measurement sets is both costly and time‐consuming for MNOs [7], which poses a significant barrier to widespread deployment.
II-B Minimization of Drive Tests (MDT)
To enable data-driven optimization, we leverage the MDT framework introduced in 3GPP Release 10 [4]. In MDT, user equipments periodically report field measurements, including RSRP, into the network’s operation & maintenance (O&M) system in order to facilitate failure forecasting, troubleshooting, and network optimization [36, 30]. While this framework provides a rich, cost-efficient alternative to traditional drive testing, it is subject to the operational constraints of commercial user devices. Notably, in the commercial MDT dataset analyzed in this work, valid GPS coordinates are present in fewer than 4% of the reported samples, as devices frequently disable satellite positioning to conserve power. This sparsity necessitates the development of augmentation strategies that can effectively utilize the limited available geolocated data.
III Related Work
Data augmentation has emerged as a powerful technique across several domains, including image classification [20, 24] and audio recognition [35], to mitigate data scarcity, enhance generalization, and improve model robustness. In the context of wireless positioning, these challenges are even more pronounced due to the high cost and manual effort involved in collecting ground-truth fingerprints in real-world environments, motivating the development of augmentation strategies that can synthetically expand datasets while preserving the spatial and statistical characteristics of the underlying radio environment. While established geo-statistical estimators like Kriging explicitly model spatial correlation to produce geo-tagged features at unmeasured locations [28], these approaches often have high sensitivity to model assumptions, rely on the assumption of data stationarity and isotropy, and become computationally intensive for very large datasets.
A recent survey [11] has documented various augmentation strategies developed for WiFi-based indoor fingerprinting, such as spatial interpolation and signal transformations by leveraging generative models. However, these approaches are proven to be effective in dense indoor settings using short-range signal features, while their efficacy in outdoor cellular scenarios, which involve sparse, noisy, and heterogeneous measurements, is not thoroughly investigated yet. Transferring these methods to the outdoor domain is non-trivial due to distinct operational characteristics: unlike static indoor layouts, wide-area cellular networks exhibit significant variability in macro-cell density and coverage geometry, alongside practical anomalies such as incomplete PCI visibility and unpredictable user mobility patterns [23].
Several positioning approaches, which typically rely on signal propagation models [26, 46] or fingerprinting systems supported by specialized hardware deployments [18, 27], can yield accurate results in controlled environments, but they are difficult to scale in large outdoor areas where exhaustive site surveys are impractical and radio propagation is highly variable in space and time. Furthermore, operational outdoor networks are subject to continuous temporal drift driven by environmental non-stationarity (e.g., seasonal changes), traffic-dependent interference, and long-term network evolution, all of which degrade the longevity and reliability of static fingerprint databases [12]. As such, there remains a pressing need for augmentation methods tailored specifically to outdoor cellular positioning challenges. Building on this indoor-focused literature, GPR has been explored for augmenting WiFi fingerprints in [39] to improve positioning performance. Although initially designed for indoor settings, we nonetheless treat GPR as a meaningful baseline for radio-feature augmentation and include it in our experimental comparisons to quantify its strengths and limitations on outdoor data.
It is worth noting that radio features can also be predicted using deterministic approaches, such as ray tracing (RT) [15, 41], or empirical propagation models [48]. While RT can provide accurate channel estimation, it requires precise 3D environmental models (digital twins) and detailed infrastructure metadata (e.g., antenna patterns, material properties), which are computationally expensive to process and rarely available for large-scale operational networks. Similarly, standard empirical models often lack the granularity to capture local shadowing effects essential for fingerprinting. Furthermore, both approaches typically model the physical channel rather than the reported measurement data, without accounting for device heterogeneity, body-loss, and measurement noise present in real MDT traces. In contrast, data-driven approaches allow for feature augmentation that implicitly learns these complex statistical distributions directly from the data.
In a recent study, the GUMBLE framework [37] has been proposed as a general-purpose generative model for mobile network data using Bayesian conditional inference. While effective in learning conditional relationships in network data, this framework is not tailored to generate synthetic data to improve positioning accuracy; instead, it focuses on quantifying the aleatoric and epistemic uncertainty of a synthetically generated mobile dataset. Other recent generative frameworks have explored conditioning the signal generation process on auxiliary environmental features such as terrain, cell orientation, and transmission characteristics [38], while other approaches have leveraged satellite imagery or GPS traces to estimate signal quality at unmeasured locations [40]. However, such contextual inputs are rarely available in practice. Recent literature supports the view that complex generative models may not be optimal for all scenarios. For instance, the survey in [31] indicates that for low-dimensional, non-smooth radio data where network geometry is unknown, interpolation-based approaches can offer superior robustness compared to deep learning methods. Building on this insight, and in contrast with the existing literature, the focus of our work is to develop a data-driven, context-agnostic augmentation framework specifically tailored for positioning in large-scale outdoor settings. Our proposed method requires only geo-tagged datasets without auxiliary environmental metadata, and enables the generation of synthetic user locations and radio features in a manner that reflects the true characteristics of real-world mobile datasets to enhance positioning accuracy, particularly in urban regions where data scarcity and signal irregularities are most pronounced.
Longitude Latitude RSRP_PCI_1 RSRP_PCI_2 RSRP_PCI_3 $\dots$
11.3456 44.4945 -87 dBm -95 dBm -- $\dots$
11.3460 44.4951 -90 dBm -92 dBm -105 dBm $\dots$
11.3465 44.4958 -85 dBm -- -102 dBm $\dots$
11.3470 44.4963 -88 dBm -96 dBm -113 dBm $\dots$
$\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$ $\ddots$
Table I: Representative sample of the MDT dataset.
IV Mobile Data Augmentation
Figure 1: Mobile data augmentation architecture, comprising Spatial Features Augmentation and Radio Features Augmentation stages. The augmented dataset is ready to use for fingerprinting-based positioning algorithms.
The goal of the proposed mobile data augmentation architecture is to synthesize large volumes of geographically coherent, multi-cell radio features that (i) fill spatial gaps in operator-collected traces and (ii) preserve the statistical characteristics of real measurements, so that the augmented data can be used directly by fingerprinting-based positioning algorithms. The architecture is intentionally computationally light and comprises two modular stages: Spatial Features Augmentation and Radio Features Augmentation. Fingerprints built from geo-referenced per-cell power measurements have been widely used for positioning across data sources, ranging LTE-RSRP collected in field trials and drive tests [6, 17] to scanner-based datasets [7]. While broadly applicable across mobile-network data sources, we focus on MDT as a representative case due to its 3GPP compliance and widespread adoption by MNOs [4]. However, it is important to note that our approach is not specific to MDT and can be applied to any geo-tagged dataset with per-cell power measurements, such as those collected through mobile network scanners, drive tests, or other crowdsourced data campaigns.
Starting from a set of geo-tagged MDT records collected by the MNO, the proposed pipeline proceeds in two stages. First, KDE is used to estimate the empirical spatial distribution of measurements and to draw synthetic sample locations. Second, for each synthetic location, a multi-cell RSRP vector is predicted via a KNN operating on the original MDT fingerprints. This produces augmented MDT data that can be integrated into fingerprinting-based positioning algorithms. The overall flow is illustrated in Fig. 1. MDT data structure is discussed in Sec. IV-A, while the two-stage architecture details are given in Sec. IV-B and Sec. IV-C.
IV-A MDT Data
This study leverages a real-world dataset collected in MDT format by a commercial Italian MNO. The core radio feature in MDT data utilized for positioning is the RSRP, which quantifies the received power of the long term evolution (LTE) cell-specific reference signal (CRS) as measured at the user equipment (UE). RSRP is particularly suitable for positioning applications due to its insensitivity to co-channel interference and average network load. This robustness arises from the UEs ability to isolate and correlate the CRSs, thereby extracting PCI-specific signal strength even in complex radio conditions.
Formally, the RSRP at a given location is defined as the average power measured over all resource elements that carry CRSs across the system bandwidth [2]:
$\text{RSRP}=\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{14}P_{\text{RE},ik},$ (1)
where $N$ is the number of subcarriers and $P_{\text{RE},ik}$ denotes the received power for subcarrier $i$ and orthogonal frequency division multiplexing (OFDM) symbol $k$ for each time transmission interval (TTI). The value of $P_{\text{RE},ik}$ is determined as:
$P_{\text{RE},ik}=\begin{cases}P_{\text{RE},i}&\text{if the }k\text{-th OFDM symbol carries CRS},\\ 0&\text{otherwise}.\end{cases}$ (2)
RSRP serves as a critical input to multiple radio resource management (RRM) procedures, including mobility control, cell reselection, and handover decisions. As a spatially informative and locally averaged indicator, it provides a stable reference for positioning tasks, particularly when collected across multiple cells.
In the dataset, RSRP measurements are reported not only for the serving PCI, but also for all detectable neighboring PCIs within the reception range of the corresponding UE. This structure, exemplified in Tab. I, provides a direct representation of the measurement format. Each entry is associated with a geographical position expressed as a latitude-longitude pair and RSRP measurements for each distinct PCI observed within the considered region. For positions where a given PCI is not detected, the corresponding measurement is absent, resulting in a sparse data structure that reflects realistic coverage conditions.
IV-B Spatial Features Augmentation via Density Estimation
Tabular data generation, including spatial features augmentation, can be addressed using a wide range of approaches [29]. State-of-the-art methods include GANs [33, 9], NFs [34], variational autoencoders [45], and GMMs [25]. However, the spatial distribution of MDT data presents specific challenges for these models due to its inherent complexity, irregularity, and strong dependence on the topography of the area under investigation.
To this end, we propose modeling user sample generation through a bi-variate density estimation approach over the latitude and longitude components of MDT data. Specifically, we can set the density estimation problem as a maximum log-likelihood problem:
$\theta^{*}=\operatorname*{argmin}_{\theta}\ \mathbb{E}_{\mathbf{p}\sim P(\mathbf{p})}\bigl[-\log f(\mathbf{p}\mid\theta)\bigr]\,$, (3)
where $\mathbf{p}\in\mathbb{R}^{2}$ denotes the latitude and longitude of a data point sampled from the true distribution $P(\mathbf{p})$, $\theta$ is the model parameterization, and $f(\mathbf{p}\mid\theta)$ is the modeled density. In practice, we approximate the expectation via empirical risk minimization (ERM) over a training set $\mathcal{D}=\{\mathbf{p}_{i}\}_{i=1}^{m}$, yielding the empirical loss
$\mathcal{L}(\theta)=-\sum_{i=1}^{m}\log f(\mathbf{p}_{i}\mid\theta)\,$, (4)
and optimize $\theta^{*}=\operatorname*{argmin}_{\theta}\mathcal{L}(\theta)$.
Specifically, we adopt KDE, a non-parametric technique, motivated by its simplicity, interpretability, and suitability for modeling the spatial distribution of users. Unlike black-box generative models, KDE offers a transparent decision mechanism where the estimated density is directly traceable to the proximity of specific training examples. In Sec. VI-A, we further validate this choice by comparing performance using the multivariate KS test against parametric baselines such as GANs, NFs, and GMMs. Specifically, KDE estimates the underlying probability density function (PDF) by placing a kernel $K(\mathbf{p}-\mathbf{p}_{i},h)$ centered at each training point $\mathbf{p}_{i}\in\mathcal{D}$, with $h$ representing the kernel bandwidth. This yields an interpretable PDF estimate $f(\mathbf{p},\mathcal{D},h):\mathbb{R}^{2}\rightarrow\mathbb{R}^{+}$, constructed directly from the empirical data, capturing multi-modal structures inherent in the distribution [44]. Formally, the KDE is defined as
$f(\mathbf{p},\mathcal{D},h)=\frac{1}{m}\sum\limits_{i=1}^{m}K(\mathbf{p}-\mathbf{p}_{i},h)\,$, (5)
where the kernel is given by the Gaussian function:
$K(\mathbf{p}-\mathbf{p}_{i},h)=\frac{1}{h\sqrt{2\pi}}\exp\left(-\frac{(\mathbf{p}-\mathbf{p}_{i})^{2}}{h^{2}}\right)\,.$ (6)
Here, $\mathbf{p}$ is the evaluation point, and the model parameters $\theta$ correspond to the training data points $\mathcal{D}$.
We select the Gaussian kernel based on domain-specific considerations. MDT location data are typically obtained via GPS at the user side, and the associated measurement noise is typically modeled as Gaussian. Therefore, using a Gaussian kernel with a tunable bandwidth aligns naturally with the assumed uncertainty in user positioning. This reinforces the model’s interpretability: the predicted density at any point $\mathbf{p}$ is not an abstract score, but a cumulative measure of how many training users were physically located within the noise radius $h$ of that point. Intuitively, KDE assumes that regions with higher sample density are more likely to generate future observations. Since KDE is an unsupervised method, the bandwidth $h$ is treated as a hyperparameter and tuned via ERM. Specifically, we minimize the negative log-likelihood on a validation subset $\mathcal{D}^{\prime}=\{\mathbf{p}_{i}^{\prime}\}_{i=1}^{m^{\prime}}$, that is
$\operatorname*{argmin}_{h}-\sum_{i=1}^{m^{\prime}}\log f(\mathbf{p}_{i}^{\prime},\mathcal{D},h)\,.$ (7)
IV-C Radio Features Augmentation via Non-Parametric Regression
Radio features augmentation can be addressed with a range of approaches; representative examples include neural networks [37], RFs [42], and GPRs [39], which have been used for related tasks in literature. In this work, we adopt KNN, which is a simple, interpretable non-parametric procedure that transfers measured per-PCI radio features from nearby reference locations. In Sec. VI-B, we further validate this choice by comparing performance using MedAE on radio feature prediction against machine learning approaches (RF and GPR) and deep learning approaches (MLP, CGAN, ReFlow, and NF).
Let $N_{c}$ denote the total number of cells in the region of interest. We assume access to a reference fingerprint database
$\hat{\mathbf{R}}=\bigl\{(\hat{\mathbf{r}}_{j},\;\hat{\mathbf{p}}_{j})\bigr\}_{j=1}^{m}\quad\text{with}\quad\hat{\mathbf{r}}_{j}\in\mathbb{R}^{N_{c}},\;\hat{\mathbf{p}}_{j}\in\mathbb{R}^{2},$ (8)
where each $\hat{\mathbf{r}}_{j}$ is a measurement vector recorded at the known location $\hat{\mathbf{p}}_{j}$. Given a query spatial location $\mathbf{p}\in\mathbb{R}^{2}$ (either sampled from a test set, sampled from the KDE in Sec. IV-B, or otherwise selected), our goal is to synthesize a realistic multi-cell radio fingerprint ${\mathbf{r}}(\mathbf{p})=[r_{1}(\mathbf{p}),\dots,r_{N_{c}}(\mathbf{p})]^{\top}$ consisting of per-PCI RSRP measurements that are consistent with the empirical database $\hat{\mathbf{R}}$.
We use a spatial nearest-neighbor rule to transfer radio measurements from the reference database to the query point, where the measurements of the closest reference point are assigned to the query point. We denote the index of the training location closest in Euclidean distance to the query point $\mathbf{p}$ as
$j^{*}(\mathbf{p})\;=\;\operatorname*{argmin}_{j=1,\dots,m}\bigl\lVert\mathbf{p}-\hat{\mathbf{p}}_{j}\bigr\rVert_{2}\,$, (9)
where $\bigl\lVert\cdot\bigr\rVert_{2}$ denotes the Euclidean distance. The nearest training fingerprint $\hat{\mathbf{r}}_{j^{*}}$ provides the baseline per-PCI measurements to be assigned to $\mathbf{p}$. To model slow, large-scale fluctuations caused by obstacles, buildings and other shadowing phenomena, we perturb the assigned baseline RSRP values using a log-normal (additive Gaussian in the dB scale) shadowing term, which is the standard parametrization in the propagation literature [3]. Concretely, for each PCI $c$, we draw an independent shadowing sample
$s_{c}\;\sim\;\mathcal{N}\bigl(0,\sigma_{s}^{2}\bigr),$ (10)
and define the augmented (continuous) measurement
$r_{c}(\mathbf{p})=\hat{r}_{j^{*},c}+s_{c}\,.$ (11)
Notably, if a PCI is not detected in the $j$-th training location, we simply leave it unassigned, i.e., no measurement is generated for that PCI at $\mathbf{p}$.
IV-D Analysis and Benefits
The proposed mobile data augmentation pipeline is intentionally simple and modular: a centralized Spatial Features Augmentation stage (KDE-based) produces geographically coherent locations, while a lightweight Radio Features Augmentation stage (KNN-based) transfers and perturbs empirical per-PCI RSRP fingerprints. This separation of concerns is the starting point of several practical advantages that make the architecture attractive for operator-grade deployment.
  1. 1.
    The architecture naturally supports a hybrid deployment model. The spatial KDE, requiring a global view of geo-tagged reports to accurately capture multi-modal spatial densities, is best executed centrally (e.g., within the operator back-end or an on-premise data-hub). Conversely, the radio-transfer KNN-based module can be executed in a distributed manner (e.g., at the network edge, in regional data-centers, or even on-site at measurement collection points) because it only needs access to a local fingerprint catalogue or indexed subset of $\hat{\mathbf{R}}$. This division reduces communication overhead: only the spatial model or sampled synthetic locations must be exchanged, while radio fingerprints can be generated locally when required.
  2. 2.
    The pipeline is essentially training-free for radio-feature synthesis. KDE sampling and the nearest-neighbor algorithm avoid iterative, parameter-intensive training loops typical of RF, GPR, or deep models. The practical consequences are significant: lower computational and engineering costs, faster time-to-deployment, and elimination of long retraining cycles whenever new MDT records become available. Because no complex model parameters must be learned and propagated, newly collected data can be appended and used immediately to refresh the empirical database or to re-sample the KDE, providing a seamless data-integration path that directly benefits both augmentation and downstream positioning.
  3. 3.
    The approach yields high interpretability and simplifies validation. KDE bandwidth, kernel choice, and the KNN algorithm are explicit, easy-to-inspect hyperparameters whose effects on the generated distribution are intuitive. This transparency facilitates validation and supports operator certification workflows, where explainability and repeatable test procedures are often mandatory. In practice, this means operators can reliably audit the augmentation process, identify failure modes, and enforce quality gates before synthetic data are admitted into production positioning databases.
  4. 4.
    The design supports privacy- and data-minimization strategies. Since the spatial model can be represented by a compact density parameterization (or by synthetic locations sampled from it), operators can avoid sharing raw, user-level MDT traces across different sites. Moreover, in distributed deployments, the detailed per-PCI measurements can remain in the data domain where they were collected, reducing privacy exposure and simplifying compliance with data-protection requirements.
  5. 5.
    The modular architecture is extensible: the spatial features augmentation and the radio features augmentation components can be replaced or augmented with other predictors when computational resources, data volume, or fidelity requirements justify it. Because modules expose clear interfaces (i.e., the input/output scheme at each component), alternative models can be integrated or benchmarked without redesigning the whole generator or the downstream positioning pipeline.
  6. 6.
    From an operational standpoint, the combination of low runtime cost and immediate usability of newly collected data makes the KDE-KNN solution particularly suitable for on-premise or edge deployments where compute resources are constrained and timeliness is important.
Taken together, these properties explain why a lightweight, statistically faithful augmentation pipeline can be a practical and robust choice for improving fingerprinting-based positioning in real-world MDT-driven scenarios.
We note that the spatial- and radio-feature augmentation modules admit independent design choices and hyperparameters. For example, the KDE-based spatial stage can employ different kernel families and bandwidth-selection rules, while the radio stage may use alternative nearest-neighbor schemes (e.g., $K>1$, distance-weighted averaging, or diverse distance metrics) as well as varying shadowing and noise models. A systematic exploration of these alternatives and their impact on downstream positioning performance is deferred to future work.
V Fingerprinting-based Positioning and
Real-World MDT-based Scenarios
V-A Multi-cell Fingerprinting-based Positioning
There exist several neural network-based approaches to multi-cell fingerprinting-based positioning, including those in [14, 22]. Other works, including [16, 7, 8], rely on non-parametric predictors such as weighted k‐nearest‐neighbor (wKNN) for determining position information. In this work, we leverage wKNN as an illustrative example of a positioning algorithm that can possibly benefit from mobile data augmentation.
In multi-cell fingerprinting-based positioning, a set of base stations (cells) act as anchors whose received‐signal measurements are used to infer an unknown user location. As introduced in Sec. IV-C, we denote by $\hat{\mathbf{R}}$ the reference fingerprint database, comprising $m$ radio measurements with respect to $N_{c}$ anchors and user position information. For the ii‑th user (at unknown position $\mathbf{p}_{i}\in\mathbb{R}^{2}$), we observe a radio‑measurement vector
$\mathbf{r}_{i}=\bigl[r_{i,1},\,r_{i,2},\,\dots,\,r_{i,N_{c}}\bigr]^{\top}\;\in\;\mathbb{R}^{N_{c}}\,.$ (12)
To estimate $\mathbf{p}_{i}$ from $\mathbf{r}_{i}$, we employ the wKNN algorithm. For any two measurement vectors $\mathbf{r}_{i},\hat{\mathbf{r}}_{j}\in\mathbb{R}^{N_{c}}$, we define a distance
$d(\mathbf{r}_{i},\hat{\mathbf{r}}_{j})=\bigl\lvert\mathbf{r}_{i}-\hat{\mathbf{r}}_{j}\bigr\rvert={\sum_{c=1}^{N_{c}}\;\bigl\lvert\mathbf{r}_{i,c}-\hat{\mathbf{r}}_{j,c}\bigr\rvert}\,$, (13)
and we perform neighbor selection computing all distances $d\bigl(\mathbf{r}_{i},\hat{\mathbf{r}}_{j}\bigr)$ for all $j=1,\dots,m$. We denote the indices of the $K$ reference fingerprints with the smallest distance to $\mathbf{r}_{i}$ as
$\mathcal{N}_{K}(i)\;=\;\bigl\{\,j_{1},\dots,j_{K}\bigr\}\subset\{1,\dots,m\}\,.$ (14)
Then, we assign to each neighbor $j\in\mathcal{N}_{K}(i)$ a weight
$w_{j}=w\bigl(d(\mathbf{r}_{i},\hat{\mathbf{r}}_{j})\bigr)=\frac{1}{\,d(\mathbf{r}_{i},\hat{\mathbf{r}}_{j})+\varepsilon\,}\,$, (15)
where $\varepsilon\in\mathbb{R}^{+}$ is a small positive constant. The estimated position $\mathbf{p}_{i}$ is the normalized weighted average of the true positions of the $K$ nearest neighbors, that is
$\mathbf{p}_{i}=\frac{\displaystyle\sum_{j\in\mathcal{N}_{K}(i)}w_{j}\,\hat{\mathbf{p}}_{j}}{\displaystyle\sum_{j\in\mathcal{N}_{K}(i)}w_{j}}\,.$ (16)
This wKNN framework provides a simple, non‑parametric baseline for fingerprinting‑based positioning. In the sequel, we will assess how augmenting the reference database $\hat{\mathbf{R}}$ can improve the accuracy of this estimator.
V-B MDT-based Reference Scenarios
In this work, we leverage a real-world MDT dataset collected across the city of Bologna (Italy) and provided by an Italian MNO. In the dataset, each sample corresponds to a user record, including its latitude, longitude, and RSRP measurements that are reported for all visible PCIs within its reception range. To ensure the reliability of these spatial labels subject to GPS multipath and drift, we applied a quality control filter based on the positioning uncertainty fields defined in 3GPP standards. Specifically, we utilized the uncertainty semi-major axis to discard samples with a positioning uncertainty radius exceeding 10 m. This pre-processing step effectively mitigates the impact of location drift, ensuring that the ground-truth coordinates used for training remain robust even in challenging scenarios. The measurements cover four distinct regions, namely a dense urban area, mixed urban-rural zones, an airport, and a highway. To study the impact of augmentation under diverse environmental and operational conditions, the four scenarios are chosen to be heterogeneous.
These regions vary in spatial density, mobility profiles, and radio coverage characteristics, offering a representative basis for evaluating RSRP-based fingerprinting positioning systems. To reflect realistic propagation environments, each area is modeled under predominantly non-line-of-sight (NLOS) conditions, which are common in urban and suburban deployments. Additionally, a shadow fading term is introduced for each region following the 3GPP Release 10 [4] standard and incorporated into the PCI-wise radio feature modeling process. Visual representations of these regions are provided in Fig. 2. Their quantitative characteristics, including spatial extent, number of UEs, user density, and shadow fading variance $(\sigma_{s}^{2})$, are summarized in Tab. II.
(a) City center
(b) Stadium
(c) Airport
(d) Highway
Figure 2: The four MDT-based regions in Bologna representing our reference scenarios.
Region Area (km2) Num. of UEs Density (UE/km2) $\bm{\sigma^{2}_{s}}$ (dB)
City center 2.05 6453 4190 8.8
Stadium 3.20 2705 1253 7.8
Airport 18.61 17698 720 7.8
Highway 24.96 2920 90 8.0
Table II: Quantitative descriptors and shadow fading variance of the four MDT-based regions.
(a) Ground-Truth
(b) KDE
(c) GMM
(d) GAN
(e) NF
Figure 3: Illustrative comparison of real MDT geo-located samples and synthetic locations generated by KDE, GMM, GAN, and NF for the airport scenario.
Model City center Stadium Airport Highway
Test statistic $p$-value Test statistic $p$-value Test statistic $p$-value Test statistic $p$-value
[…]
GAN $062\pm 0.025 0.0$$50\pm 0.086 0.0$$96\pm 0.028 0.0$$45\pm 0.078 0.0$$32\pm 0.013 0.0$$69\pm 0.086 0.0$$064\pm 0.023$ $0.300\pm 0.262$
NF $0.029\pm 0.004$ $277\pm 0.104 0.0$$040\pm 0.008$ $348\pm 0.192 0.0$$25\pm 0.004 0.0$$61\pm 0.047 0.0$$048\pm 0.011$ $0.374\pm 0.209$
Table III: Multivariate two-sample KS test statistics and $p$-values comparing synthetic spatial samples generated by KDE, GMM, GAN and NF against real MDT locations for each reference region. The null hypothesis (real and synthetic samples drawn from the same distribution) is rejected for $p\text{-value}<0.05$.
The selected areas include:
  • The city center area, which represents the most compact and dense region and covers the historical core of Bologna and is served by 140 distinct PCIs.
  • The stadium area, which corresponds to a mixed urban-rural zone covered by 104 PCIs and comprises a dense cellular deployment outside the historical core of the city.
  • The airport area, which is the largest zone and exhibits moderate user mobility, with service provided by 209 PCIs. While it contains the second-highest number of UEs after the city center, its users are distributed across a wider area, resulting in a reduced user density compared to the city center and stadium areas.
  • The highway area, which is located in the northern part of Bologna and features low user density and high user mobility, with service provided by 52 PCIs. It represents a large-scale vehicular scenario.
VI Numerical Results
This section evaluates the proposed mobile data augmentation architecture both stage-by-stage and end-to-end, emphasizing its effect on fingerprinting-based positioning. We first validate each module in isolation to verify that the Spatial Features Augmentation and the Radio Features Augmentation behave as intended (Sec. VI-A and Sec. VI-B, respectively). Then, we assess the impact of the full augmentation pipeline on multi-cell fingerprinting-based positioning across the four real-world MDT scenarios introduced in Sec. V-B (Sec. VI-C).
Our primary objective is to quantify improvements in positioning performance, while also ensuring that intermediate modules produce statistically faithful and useful synthetic data. Accordingly, module evaluations report relevant per-module metrics (i.e., multivariate two-sample KS tests and radio prediction error metrics such as MedAE), whereas the end-to-end study focuses on positioning accuracy. All results compare against established baselines for the respective tasks, while Sec. VI-D investigates the scaling effect of progressively augmenting the reference database (i.e., how adding more synthetic samples affects positioning).
VI-A Results on Spatial Features Augmentation
We compare four approaches for modeling the empirical spatial distribution of MDT samples [37]: (i) the proposed Gaussian KDE; (ii) a GMM fitted via expectation-maximization with full covariances; (iii) a GAN with fully-connected generator and discriminator trained adversarially with Adam; and (iv) a likelihood-based NF trained by maximum likelihood. All methods are trained on the same region-specific training sets and generate synthetic 2D spatial locations.
To assess whether synthetic samples are statistically similar to real MDT locations we employ a multivariate two-sample KS test. Using this test, we assess whether synthetic user positions generated by KDE and the competing methods are statistically consistent with the spatial distribution of positions in the original test set. Under the null hypothesis, the real samples and synthetic samples originate from the same distribution; we reject the null whenever the KS $p$-value falls below the significance level $\alpha=0.05$. Tab. III reports the average test statistic and the average $p$-value for each model and region over ten independent runs, with associated 95% confidence interval (CI). Fig. 3 provides an illustrative example of real and synthetic samples for one representative region.
Overall, KDE and GMM consistently produce the smallest KS statistics and yield $p$-values well above $\alpha$ across all four regions, indicating no statistically significant difference from the empirical spatial distribution, while GAN and NF approaches may require additional regularization or architecture tuning to match the empirical MDT distributions in the most heterogeneous regions. These results validate the use of KDE, chosen for its interpretability and light computational footprint, as a reliable, lightweight method for spatial sample generation in our mobile data augmentation architecture.
VI-B Results on Radio Features Augmentation
Similarly, for radio feature augmentation we compare eight methods for synthesizing per-PCI RSRP fingerprints [39]: (i) the proposed spatial KNN-based augmentation with additive log-normal shadowing; (ii) a RF regression trained per-PCI; (iii) a GPR with a SE kernel; and (iv) a GPR with a RQ kernel; (v) an MLP with three hidden layers; (vi) a CGAN; (vii) a ReFlow model, representing state-of-the-art diffusion-based generation; and (viii) a NF model. All radio feature models are trained on region-specific training sets and evaluated on held-out test sets using the MedAE on per-PCI RSRP values. This evaluation is conducted independently for each geographic region. The reported values in Tab. IV represent the average MedAE over ten independent runs, with associated 95% CI.
The results demonstrate that all models achieve relatively low MedAE values, indicating effective RSRP estimation across diverse urban and semi-urban regions. The KNN model consistently yields the lowest MedAE across all regions, particularly excelling in the airport and city center zones, which contain the highest number of UEs. In contrast, the GPR models show slightly worse performance overall. The RQ variant consistently outperforms the standard GPR across all regions, as it incorporates a more flexible kernel capable of modeling multi-scale signal variations. RF exhibits moderate performance, with accuracy generally falling between KNN and GPR. Moreover, the experimental results reveal a relevant trend: lightweight, instance-based methods (specifically KNN) consistently outperform complex deep generative architectures (such as CGAN, ReFlow, and NF) in RSRP estimation. This performance gap is primarily driven by the context-agnostic nature of the problem, where the input space is restricted to low-dimensional spatial coordinates (latitude, longitude) without auxiliary environmental metadata. While deep learning models typically excel at extracting hierarchical features from high-dimensional structured data, they often struggle to model the high-frequency spatial variations inherent to radio propagation (e.g., local shadowing and multipath) when constrained to such sparse inputs, which is a phenomenon often linked to the spectral bias of neural networks [32]. In contrast, KNN operates as a non-parametric learner with a strong local inductive bias. By avoiding assumptions about global function smoothness, it effectively captures the localized and irregular nature of the radio environment, proving that in the absence of geometric features, local interpolation remains superior to deep functional approximation.
Model City center Stadium Airport Highway
[…]
Table IV: MedAE of RSRP prediction for considered approaches in each reference region.
VI-C Results on Multi-cell Fingerprinting-based Positioning
In this section, we evaluate the end-to-end impact of the proposed mobile data augmentation architecture on multi-cell fingerprinting-based positioning. Starting from the region-specific reference database, we compare positioning performance obtained with the original operator-collected fingerprints against databases augmented with synthetic samples produced by the spatial and radio augmentation stages. The positioning algorithm under test is the wKNN estimator introduced in Sec. V-A; results are reported for all four MDT reference scenarios.
We compare our proposed two-stage architecture KDE-KNN against seven alternative two-stage configurations: KDE-RF, KDE-GPR (SE/RQ), KDE-MLP, KDE-CGAN, KDE-ReFlow, and KDE-NF. For each MDT scenario, we evaluate a set of augmentation rates $A\in\{1,5,10,20,30\}$, where $A$ denotes the multiplicative increase of the fingerprinting database relative to the original ($A=1$ corresponds to no augmentation). For every combination of scenario, augmentation rate and architecture, we perform ten independent runs. In the remanding of this section, we report average MedAE values for positioning with 95% CI.
  1. 1.
    City Center Area (Tab. V). In the city center scenario, which features the highest measurement density and the smallest spatial extent, the wKNN estimator applied to the original MDT baseline attains a median positioning error of 20.0820.08\,m. With the proposed KDE-KNN pipeline and an augmentation factor $A=20$, the median error decreases to 17.8017.80\,m, corresponding to an $\approx 11.4\%\$relative improvement. Although the absolute reduction ($\approx 2.3\,$m) is modest, it is noteworthy given the strong baseline performance. Moreover, KDE-KNN consistently outperforms the alternative two-stage configurations, which yield smaller or no gains; in particular, the KDE-RF variant may be detrimental, since RF-based RSRP predictions can produce augmented fingerprints that degrade positioning accuracy.
  2. 2.
    Stadium Area (Tab. VI).
    Augm. rate $A$ KDE-KNN KDE-RF KDE-GPR KDE-GPR(RQ) KDE-MLP KDE-CGAN KDE-ReFlow KDE-NF
[…]
    Table V: Impact of data augmentation rates on positioning across increasing synthetic data ratios in the city center scenario.
    Augm. rate $A$ KDE-KNN KDE-RF KDE-GPR KDE-GPR(RQ) KDE-MLP KDE-CGAN KDE-ReFlow KDE-NF
[…]
    Table VI: Impact of data augmentation rates on positioning across increasing synthetic data ratios in the stadium scenario.
    In the stadium scenario, characterized by a high measurement density but a very confined spatial extent and an intermediate number of PCIs, the proposed KDE-KNN pipeline yields substantial gains. With augmentation factor $A=20$, the median positioning error drops from $36.70$ m to $28.32$ m, an absolute improvement of $8.38$ m and a relative reduction of $\approx 23\%$. By contrast, the alternative two-stage configurations provide only marginal or no improvements across augmentation levels.
  3. 3.
    Airport Area (Tab. VII).
    Augm. rate $A$ KDE-KNN KDE-RF KDE-GPR KDE-GPR(RQ) KDE-MLP KDE-CGAN KDE-ReFlow KDE-NF
[…]
    Table VII: Impact of data augmentation rates on positioning across increasing synthetic data ratios in the airport scenario.
    In the airport scenario, which corresponds to the largest area in our study and the one served by the highest number of PCIs, the propagation environment and spatial variability are markedly more complex. Under these conditions, the KDE-KNN pipeline yields only a modest improvement ($\approx$4%) at $A=20$, while other alternative augmentation schemes actually degrade positioning performance. These findings underline that simply increasing the quantity of synthetic samples is not sufficient: augmentation must preserve the statistical and spatial structure of the radio field, otherwise imperfect radio predictors can introduce distortions that harm positioning accuracy.
  4. 4.
    Highway Area (Tab. VIII).
    Augm. rate $A$ KDE-KNN KDE-RF KDE-GPR KDE-GPR(RQ) KDE-MLP KDE-CGAN KDE-ReFlow KDE-NF
[…]
    Table VIII: Impact of data augmentation rates on positioning across increasing synthetic data ratios in the highway scenario.
    In the highway scenario, where UE measurements are highly concentrated along the linear roadway (i.e., samples cluster on the highway) but sparse elsewhere across the largest geographic extent and subject to high mobility, augmentation delivers the largest benefit. With the proposed KDE-KNN pipeline and augmentation factor $A=30$, the median positioning error falls from $115.51$ m to $79.82$ m, a relative reduction of $\approx 30\%$. This substantial improvement indicates that realistic augmentation can compensate for the anisotropic and sparse spatial sampling typical of highway measurements; by contrast, the alternative two-stage configurations produce markedly smaller gains and are consistently outperformed by KDE-KNN across augmentation levels.
These findings lead to three relevant conclusions: (i) the KDE-KNN pipeline consistently improves positioning accuracy across all evaluated scenarios, with the largest relative gains in spatially sparse or structurally complex regions; (ii) the alternative two-stage baselines fail to match the proposed approach across the tested augmentation rates; (iii) performance gains saturate beyond a region-specific augmentation factor, so adding more synthetic samples does not by itself guarantee better accuracy, and poorly matched synthetic fingerprints can even degrade it; and (iv) in high-density scenarios (e.g., city center or stadium), the saturation in terms of positioning performance occurs at low augmentation rates and converges to low values, as UE density is already high in the training set, as shown in Table II. Conversely, in the highway scenario, characterized by high mobility and extremely low spatial density, the impact of data sparsity is significant. Consequently, this region requires much higher augmentation rates before saturation is reached and overall positioning algorithms perform worse.
VI-D Statistical Model Comparison
$i$ $j$ x1 x5 x10 x20 x30
[…]
x10 $cellcolorbetter \cellcolordiaggray \cellcolorneutral \cellcolorneutral$
x20 $cellcolorbetter \cellcolorneutral \cellcolordiaggray \cellcolorneutral$
x30 $cellcolorbetter \cellcolorneutral \cellcolorneutral \cellcolordiaggray$
$i$ $j$ x1 x5 x10 x20 x30
[…]
x10 $cellcolorbetter \cellcolordiaggray \cellcolorworse \cellcolorworse$
x20 $cellcolorbetter \cellcolorbetter \cellcolordiaggray \cellcolorneutral$
x30 $cellcolorbetter \cellcolorbetter \cellcolorneutral \cellcolordiaggray$
$i$ $j$ x1 x5 x10 x20 x30
[…]
x10 $cellcolorbetter \cellcolordiaggray \cellcolorworse \cellcolorneutral$
x20 $cellcolorbetter \cellcolorbetter \cellcolordiaggray \cellcolorneutral$
x30 $cellcolorbetter \cellcolorneutral \cellcolorneutral \cellcolordiaggray$
$i$ $j$ x1 x5 x10 x20 x30
[…]
x10 $cellcolorbetter \cellcolordiaggray \cellcolorneutral \cellcolorworse$
x20 $cellcolorbetter \cellcolorneutral \cellcolordiaggray \cellcolorworse$
x30 $cellcolorbetter \cellcolorbetter \cellcolorbetter \cellcolordiaggray$
X Model $i$ and Model $j$ are not statistically different X Model $i$ is statistically worse than Model $j$ X Model $i$ is statistically better than Model $j$
(a) City center             (b) Stadium             (c) Airport            (d) Highway
Figure 4: Statistical comparison of KDE-KNN model variants across four deployment scenarios. Each subfigure corresponds to a cross-model comparison at increasing augmentation rates.
The goal of this subsection is to quantify how positioning performance of the proposed KDE-KNN architecture scales with the number of synthetic fingerprints and to test whether performance gains saturate as the augmentation factor increases. In order to statistically compare model performance metrics in a pairwise manner across augmentation rates, we evaluate whether the difference $d=e_{A}-e_{B}$ between two model positioning errors $e_{A}$ and $e_{B}$ for models $A$ and $B$ is statistically significant [43, 10]. We consider two models that are evaluated $N_{e}$ times, yielding individual positioning errors $e_{A,i}$ and $e_{B,i}$ for $i=1,\dots,N_{e}$. We model the sample mean for the two compared models as
$\overline{e}_{A}=\frac{1}{N_{e}}\sum_{i=1}^{N_{e}}e_{A,i}\penalty 10000\ ,\quad\quad\overline{e}_{B}=\frac{1}{N_{e}}\sum_{i=1}^{N_{e}}e_{B,i}\penalty 10000\ ,$ (17)
and their sample variance as
$\hat{\sigma}_{A}^{2}=\frac{\sum_{i=1}^{N_{e}}\bigl(e_{A,i}-\overline{e}_{A}\bigr)^{2}}{N_{e}-1}\penalty 10000\ ,\quad\hat{\sigma}_{B}^{2}=\frac{\sum_{i=1}^{N_{e}}\bigl(e_{B,i}-\overline{e}_{B}\bigr)^{2}}{N_{e}-1}\penalty 10000\ .$ (18)
For $N_{e}\geq 30$, by the Central Limit Theorem [21], $e_{A}$ and $e_{B}$ converge in distribution to $\mathcal{N}\!\bigl(\overline{e}_{A},\;\hat{\sigma}_{A}^{2}/N_{e}\bigr)$ and $\mathcal{N}\!\bigl(\overline{e}_{B},\;\hat{\sigma}_{B}^{2}/N_{e}\bigr)$, respectively. Consequently, $d$ converges in distribution to $\mathcal{N}\!\bigl(\bar{d},\;\hat{\sigma}_{d}^{2}\bigr)$, where $\bar{d}=\overline{e}_{A}-\overline{e}_{B}$ and $\hat{\sigma}_{d}^{2}=(\hat{\sigma}_{A}^{2}+\hat{\sigma}_{B}^{2})/N_{e}$. For confidence level $1-\alpha$, the CI for $\mathbb{E}[d]$ is determined by
$\mathbb{E}[d]\in\left[\bar{d}-z_{\alpha/2}\,\frac{\hat{\sigma}_{d}^{2}}{\sqrt{N_{e}}}\,,\bar{d}+z_{\alpha/2}\,\frac{\hat{\sigma}_{d}^{2}}{\sqrt{N_{e}}}\right]\penalty 10000\ ,$ (19)
where $z_{\alpha/2}$ is the standard Normal quantile. If the CI in Eq. (19) contains zero, we conclude that the difference between the two models is not statistically significant. Otherwise, we conclude that the difference is statistically significant, i.e., that model $A$ outperforms model $B$ if the entire CI lies above zero and, conversely, that $B$ outperforms $A$ if it lies below zero.
In order to perform the pairwise comparisons, we followed the methodology described above by performing $N_{e}=30$ independent runs for each model given a region and an augmentation rate, considering a confidence level of 95% (i.e., $\alpha=0.05$). Fig. 4 summarizes these pairwise comparisons for each region: rows and columns index augmentation rates $A\in\{1,5,10,20,30\}$, and each cell indicates whether the difference in the chosen positioning metric between the row- and column-level is statistically positive, negative, or not significant, highlighting both monotonic trends and saturation points. These results reveal a clear structural difference between the considered areas. In the city center scenario, which features the highest UE density, we observe that performance saturates starting at moderate augmentation rates ($A=10$). In denser scenarios, such as stadium and airport, the transition from $A=20$ to $A=30$ yields no significant gain, suggesting that the synthetic data has effectively filled the spatial gaps within the radio map. In contrast, the highway scenario, which features the lowest UE density, exhibits a sustained improvement, with significant gains persisting up to $A=30$. This behavior is driven by the highly sparse nature of MDT data in this area. Consequently, we conclude that the saturation point for positioning accuracy is inversely related to the initial UE density of the training data: sparse topologies (e.g., highway) exhibit sustained performance gains at higher augmentation rates, whereas dense environments rapidly converge to a saturation point.
VII Conclusion
This paper has presented a practical, modular framework for mobile data augmentation tailored to real-world MDT traces collected by MNOs. The proposed two-stage architecture decouples spatial sample generation from radio-feature synthesis through KDE and KNN, delivering an interpretable and computationally lightweight solution that directly leverages operator-held geo-tagged measurements. Extensive experimental validation on four large-scale operational datasets demonstrates that the proposed pipeline yields consistent accuracy gains, particularly in scenarios characterized by sparse sampling and high mobility. Specifically, our KDE-KNN framework reduced the median positioning error by approximately $30\%$ in the sparse highway scenario and by $23\%$ in the complex stadium environment, significantly outperforming more complex deep learning and generative baselines. Furthermore, our analysis revealed a critical structural dependency between augmentation efficacy and the native density of the environment. We observed that benefits from augmentation saturate more rapidly in dense urban centers than high-mobility and sparse regions. This underscores a key operational insight: augmentation strategies must be dynamically tuned to local data sparsity, as indiscriminate over-sampling in already dense areas yields diminishing returns and may not introduce useful information. In sum, the proposed KDE-KNN augmentation offers a simple, interpretable, and deployment-ready path to more accurate outdoor fingerprinting on real MDT data, while providing operator-relevant guidance on when additional synthetic samples effectively improve positioning.